# Boosting Active Defense Persistence: A Two-Stage Defense Framework Combining Interruption and Poisoning Against Deepfake

### v1

@@ II-A Single-Function Perturbation Methods
Data poisoning is an active defense strategy that interferes with a deepfake model’s learning process by injecting perturbations into the training data. Prior works have explored various approaches that primarily target specific pre-processing steps to contaminate the training data pipeline. Some methods, such as Landmark Breaker [19], focus on a downstream task by disrupting the extraction of facial key points, which aims to degrade the final model’s quality by providing it with poorly aligned facial data for training. Other methods attack the more basic stage of the pipeline, such as Face Poison [22] and SSOD [23], which directly target the detection process itself. A failure at this initial detection stage can terminate all subsequent processing steps, such as cropping and landmark alignment, preventing the attacker from acquiring usable facial data in the first place.
@@
However, the preceding analysis shows that existing active defenses are often specialized. Interruption-based methods provide real-time protection but lack persistence against retraining, while poisoning-based methods can disrupt retraining but are ineffective against existing pre-trained models. This highlights the need for a more comprehensive framework. Our work TSDF is proposed to address this challenge by synergistically integrating these two distinct and defensive paradigms.
@@ II-B Multifunctional Perturbation Methods
To provide more comprehensive protection than single-function approaches, some research has begun to explore multifunctional methods that combine different defensive goals. A common direction has been to merge interruption with traceability. For instance, Dual Defense [25] pioneers a dual-effect defense through a single embedded watermark that synergistically provides traceability while working to disrupt face-swapping models. Similarly, Zhu et al. [26] introduce an information-containing perturbation designed to combat facial manipulation systems by embedding an identity message, thus enabling active forensics within the interrupted process. Wu et al. [27] propose a framework to resolve the common conflict between watermarking and detection. Their method fine-tunes a watermark, increasing a deepfake detector’s accuracy on watermarked images without compromising traceability.
@@
Figure 3: The overall framework of the TSDF method. The process begins in the interruption stage (upper left), where an initial perturbation ($W$) is generated by feeding a perturbed image into several deepfake models and optimizing $W$ to maximize the feature-level Mean Squared Error loss ($L_{MSE}$) and a feature enhancement loss ($L_{enh}$). Subsequently, the poisoning stage (lower left) generates a specialized perturbation to attack multiple face detectors by maximizing feature-level loss ($L_{feat}$) and output-level loss ($L_{output}$). In the combination and optimization stage (right side), the two perturbations are efficiently fused. The combined system uses a threshold $\tau$ to identify low-intensity regions in the interrupted perturbations, creating a mask where the poisoning perturbation is selectively applied.
@@ III METHOD
This section details the TSDF. As illustrated in Fig. 3, the framework is composed of three key components that work in sequence: interruption perturbation generation, poisoning perturbation injection and final fusion optimization. In the initial stage, a preliminary perturbation is designed to disrupt the feature extractors within deepfake models. This initial interruption perturbation then serves as a foundational component of the framework. It is utilized in the poisoning stage, where the perturbation targeting face detectors is precisely embedded into its low-intensity regions to create persistent protection. An optimization module fuses these two components into a unified perturbation that is visually imperceptible. The following subsections describe each component in detail.
@@ III-A Deepfake Interruption via Perturbation
Instead of directly attacking end-to-end deepfake models like CMUA [15], the proposed method targets feature extractors used in deepfake networks. This prevents feature extractors from correctly encoding facial attributes, thereby disrupting downstream attribute editing and identity reconstruction. Given an input image x, an interrupted perturbation W is introduced to generate a perturbed image $\hat{X}$ = x+W. The goal is to maximize the discrepancy between extracted feature representations.
To achieve this, we optimize the interrupted perturbation to maximize the distance between extracted features from the original and perturbed images:
@@
This loss ensures that the perturbation creates a consistent and significant distortion across the feature extractors of all targeted models. To further amplify the disruptive effect, a feature enhancement loss $\mathcal{L}_{\mathrm{enh}}$ is introduced. This component magnifies the differences between the features of the original and disturbed images by considering local, global and structural differences:
@@
Here, $\mathcal{F}_{\mathrm{enh}}$ denotes the enhanced feature. This amplification is regulated by a temperature scaling parameter $\sigma$. The term $\Delta$ represents a vector of local, global, and structural differences $(\mathrm{L_{D}},\mathrm{G_{D}},\mathrm{S_{D}})$, while $\hat{W}$ denotes the adaptive weights that control the enhancement strength. This component ensures that even small feature differences are amplified into significant disruptions, resulting in a more comprehensive and destructive feature attack.
$\mathrm{L_{D}}=Norm({E_{i}(\hat{X})})+Norm({E_{i}(x)}),$ (4)
$G_{D}=\frac{U(E_{i}(\hat{X}))-U(E_{i}(x))}{S(E_{i}(x))+e},$ (5)
$\mathrm{S_{D}}=Att(Norm({E_{i}(\hat{X})}))+Att(Norm({E_{i}(x)})),$ (6)
@@
where $Norm(\cdot)$ denotes normalization function to standardize the feature map and $U(\cdot)$ calculates the mean of a feature map across its spatial dimensions. $S(\cdot)$ calculates the standard deviation of a feature map. $Att(\cdot)$ is an attention mechanism applied to normalized feature maps to compute and weight their structural feature. $e$ is a numerical stability constant.
@@
where $\gamma$ is the step size and $\epsilon$ is the perturbation budget. The projection operator ensures that the interrupted perturbation remains within a bounded perturbation space.
@@ III-B Poisoning the Retraining Pipeline
While interruption is effective for real-term distortion of image, long-term defense must prevent an attacker from using perturbed examples to fine-tune their model. Our approach creates an effective perturbation designed to disrupt face detection, which is the crucial first step in an attacker’s data processing pipeline.
@@
$L_{\text{feat}}=\sum_{i=1}^{N}w_{i}\cdot\|F_{\text{fused}}^{i}-(-F_{\text{fused}}^{i,\text{clean}}+\epsilon\cdot M_{d})\|^{2}.$ (12)
Here, $D_{i}(\text{x})$ is the feature extracted from the face detector $i_{th}$ and $w_{i}$ is the weight of the layer $i_{th}$. $N$ is the number of feature layers to attack.
@@
$L_{total}=\mathcal{L}_{feat}+\mathcal{L}_{output}.$ (14)
@@ III-C Perturbation Fusion Optimization
The interruption perturbation and the subsequent poisoning perturbation are then synergistically combined into a unified perturbation. This fusion process is designed to ensure both defensive functions can operate cooperatively without mutual interference. To achieve this, we developed an intensity separation mechanism, the process of which is illustrated in Fig. 4.
@@
$\delta_{\text{final}}=\text{Clip}(W_{0}+\delta_{\text{t+1}},-1,1),$ (17)
@@
TABLE I: Comparison of Interruption Performance among CMUA, FOUND and TSDF. The average value is the average result of each indicator under this dataset. The results are best shown in bold.
@@
TABLE II: Poison and FDR Results of TSDF Methods.
@@ IV-A Experimental Setup
Datasets: We conduct the experiments on the CelebA dataset [31] for training. The dataset comprises over 200,000 face images, from which we use the aligned and cropped version. To maintain consistency with the baseline method FOUND [20], we randomly select 128 images as the training set for generating the perturbation. For evaluation, we then apply this generated perturbation to all images in the official test split. To further evaluate the generalization capability of TSDF, additional datasets are introduced, including the Labeled Faces in the Wild (LFW) dataset [32] and facial images extracted from the FaceForensics++ Original video dataset (FF++O) [33].
Deepfake Model: We evaluate the defense against four deepfake models, HiSD [34], AttentionGAN (AGGAN) [35], StarGAN [36], and AttGAN [37]. For StarGAN and AGGAN, we select 5 attributes: black hair, blonde hair, brown hair, gender, and age. In AttGAN’s training, the editing tasks are expanded to 14 attributes. Given HiSD’s advanced facial editing capabilities, the experiments incorporated modifications of the glasses attribute. To ensure experimental consistency and fairness, all forged models are trained on the CelebA dataset.
@@
Baseline: To comprehensively validate the framework, we establish a series of baselines from state-of-the-art (SOTA) single-function defenses. For evaluating interruption performance, our primary baselines are CMUA-Watermark [15] and FOUND [20]. For poisoning efficacy, we benchmark against Face Poison [22] and SSOD [23]. Crucially, to validate the core contribution to defense persistence, the framework’s performance after simulated retraining attack is directly contrasted with that of CMUA and FOUND. This multiple evaluation provides a rigorous validation of the framework’s superiority in both its individual functions and its overall persistence against adaptive attackers.
Implementation Details: The hyperparameter configuration is aligned with existing baseline methods, while jointly optimizing the interruption and poisoning effects of the perturbations. The perturbation range for both attacks is set to (0.05, 0.05) to balance visual fidelity with attack efficacy. In the training stage, 50 and 20 iterations are performed for interruption and poisoning, as we observed in preliminary experiments that performance saturated beyond these points. For the optimizer parameters in Equations 9 and 10, $\lambda$ and $\gamma$ are set to 0.1 and 0.001, respectively. The intensity separation threshold $\tau$ was set to 0.3, a value determined to provide the optimal balance between defensive functions, as validated in the ablation study.
@@ IV-C Compare the SOTA Methods
This section validates the effectiveness of the proposed TSDF framework through a series of direct comparisons with state-of-the-art (SOTA) methods. We begin by quantitatively demonstrating the core persistence vulnerability that compromises conventional interruption defenses when faced with a retraining attack. We then proceed to benchmark the two primary capabilities of TSDF, interruption and poisoning, against their respective SOTA counterparts as defined in the baseline selection.
Effectiveness discrepancy of Interrupted Perturbation After Retraining
To simulate a worst-case attack scenario, we create training datasets composed entirely of protected images. Specifically, we apply the SOTA interruption perturbations from CMUA and FOUND to the CelebA training set, creating two separate, fully protected datasets. The deepfake models are then retrained on these respective datasets.
The results presented in Table III reveal a critical vulnerability in conventional interruption-only methods. For instance, with the CMUA method, the average SSIM score on the CelebA dataset increases from 0.64 to 0.79 after retraining, while the average FID score falls from 209.46 to 127.42. The rising SSIM indicates that the output of perturbed examples have become more similar to the original images, while the falling FID confirms they have become more realistic and less distorted. Both metrics prove that the interruption effect has been largely nullified. This degradation provides clear quantitative evidence that retraining attacks severely compromise conventional interruption-only defenses, proving their lack of persistence. This highlights the key vulnerability our TSDF framework is designed to solve.
@@
We evaluate the interruption capability of TSDF against SOTA methods, CMUA-Watermark [15] and FOUND [20], with detailed results presented in Table I. The results show that TSDF consistently delivers a more powerful and destructive effect on the final forged images across multiple datasets and models. A key advantage of the method is its superior ability to degrade the structural integrity of the forged output. This is most obvious in SSIM scores. On the CelebA dataset, TSDF achieves an average SSIM of 0.57, which is significantly lower than the scores of CMUA and FOUND, indicating superior disruption. This trend is particularly strong against models like StarGAN, where TSDF achieves a remarkable SSIM of just 0.24, far exceeding the disruptive effect of baseline. Furthermore, this disruptive force is reflected in other key metrics. On the LFW dataset, TSDF achieves the highest average L2mask of 0.28 and the highest average FID score of 282.32, indicating the most powerful disruption among the three methods. These comprehensive results demonstrate the superior disruptive capability of TSDF compared to SOTA methods.
TABLE III: Comparison of interrupted perturbation effect before and after retraining.
@@
We then assess the effectiveness of TSDF’s poisoning component against face detectors, comparing it to random noise and reproduced versions of SSOD and FacePoison. The results in Table II are critical as they validate the core defense strategy. Instead of merely trying to survive an attacker’s adaptation, our framework fundamentally disrupts the retraining process. By significantly affecting the automated face detectors that an attacker relies on, our method makes large-scale and automated collection of training data impractical. This represents a fundamental shift in defensive strategy. The objective is no longer to create a perturbation that passively withstands retraining, but to design one that actively destroys the retraining process. This capability is crucial to achieving more persistent defense. For example, on the CelebA dataset, TSDF drastically reduces the F1-score for the RetinaFace detector to 0.78. This represents a far more successful attack compared to FacePoison and SSOD, which only achieve scores of 0.96 and 0.90 respectively. The advantage is even more pronounced against the S3FD detector, where our method achieves an F1-score of 0.72, while the other methods barely impact the detector’s performance. This superior performance is consistent across the LFW and FF++O datasets, where TSDF consistently achieves the lowest F1-score. These results validate that our framework’s poisoning component poses a significant threat to the face detection process. This capability is crucial for our two-stage defense, as it is the mechanism designed to prevent attackers from successfully retraining their models and bypassing the interruption defense.
TABLE IV: Comparative Analysis of Defense Persistence After Retraining. The table compares the interruption effectiveness of CMUA, FOUND, and the TSDF before and after a simulated retraining attack. A lower SSIM or a higher FID indicates a more effective and persistent interruption defense.
@@
TABLE V: Model Performance discrepancy After Retraining on TSDF-Poisoned Data.
@@
Figure 5: Visual comparison of the interruption effects. This figure compares the outputs of deepfake models when protected by CMUA, FOUND, and TSDF. The greater degree of visual destruction in the TSDF results show its more powerful interruption effect.
@@
TABLE VI: Experiment on input quality of datasets.
@@
IV-D Comparative Analysis of Defense Persistence
Based on the previously established vulnerability of conventional methods, we conduct a direct comparative experiment to validate the persistence of our TSDF framework. All methods are evaluated against the same adaptive retraining attack to ensure a fair comparison.
To validate the persistence of our framework, we conduct a direct comparative experiment against the SOTA interruption methods. To ensure a fair comparison, we evaluate all methods on the same two representative models, StarGAN and AttGAN. We simulated an adaptive attack by retraining these models on datasets protected by each respective defense. For TSDF, the retraining dataset was first processed by the poisoning mechanism. We then measured the interruption effectiveness of each defense before and after this retraining attack to quantify their persistence.

### v2

@@ II-A Single-Function Perturbation Methods
Data poisoning is an active defense strategy that interferes with a deepfake model’s learning process by injecting perturbations into the training data. Prior works have explored various approaches that primarily target specific pre-processing steps to contaminate the training data pipeline. Some methods, such as Landmark Breaker [19], focus on a downstream task by disrupting the extraction of facial key points, which aims to degrade the final model’s quality by providing it with poorly aligned facial data for training. However, this approach targets a downstream task that is based on the successful prior face detection. Other methods attack the more basic stage of the pipeline, such as Face Poison [22] and SSOD [23], which directly target the detection process itself. Failure in this initial detection stage can terminate all subsequent processing steps, such as cropping and landmark alignment, preventing the attacker from acquiring usable facial data in the first place.
@@
However, the preceding analysis shows that existing active defenses are often specialized. Interruption-based methods provide real-time protection but lack robustness against retraining, while poisoning-based methods can disrupt retraining but are ineffective against existing pre-trained deepfake models. This highlights the need for a more comprehensive framework. Our work TSDF is proposed to address this challenge by synergistically integrating these two distinct and defensive paradigms.
@@ II-B Multifunctional Perturbation Methods
To provide more comprehensive protection than single-function approaches, some studies have begun to explore multifunctional methods that combine different defensive goals. A common direction has been to merge interruption with traceability. For instance, Dual Defense [25] pioneers a dual-effect defense through a single embedded watermark that synergistically provides traceability while working to disrupt face-swapping models. Similarly, Zhu et al. [26] introduce an information-containing perturbation designed to combat facial manipulation systems by embedding an identity message, thus enabling active forensics within the interrupted process. Wu et al. [27] propose a framework to resolve the common conflict between watermarking and detection. Their method fine-tunes a watermark, increasing a deepfake detector’s accuracy on watermarked images without compromising traceability.
@@
Figure 3: The overall framework of the TSDF method. The process begins in the interruption stage (upper left), where an initial perturbation ($W$) is generated by feeding a perturbed image into several deepfake models and optimizing $W$ to maximize the feature-level Mean Squared Error loss ($L_{MSE}$) and a feature enhancement loss ($L_{enh}$). Subsequently, the poisoning stage (lower left) generates a specialized perturbation to attack multiple face detectors by maximizing feature-level loss ($L_{feat}$) and output-level loss ($L_{output}$). In the combination and optimization stage (right side), the two perturbations are efficiently fused. The combined system uses a threshold $\tau$ to identify low-intensity regions in the interruption perturbations, creating a mask where the poisoning perturbation is selectively applied.
@@ III METHOD
This section details the TSDF. As illustrated in Fig. 3, the framework is composed of three key components that work in sequence: interruption perturbation generation, poisoning perturbation generation and final fusion optimization. In the initial stage, a preliminary perturbation is designed to disrupt the feature extractors within deepfake models. This initial interruption perturbation then serves as a foundational component of the framework. It is utilized in the poisoning stage, where the perturbation targeting face detectors is precisely embedded into its low-intensity regions to create persistent protection. An optimization module fuses these two components into a unified perturbation that is visually imperceptible. The following subsections describe each component in detail.
@@ III-A Deepfake Interruption via Perturbation
Instead of directly attacking end-to-end deepfake models like CMUA [15], the proposed method targets feature extractors used in deepfake networks. This prevents feature extractors from correctly encoding facial attributes, thereby disrupting downstream attribute editing and identity reconstruction. Given an input image x, an interruption perturbation W is introduced to generate a perturbed image $\hat{X}$ = x+W. The goal is to maximize the discrepancy between extracted feature representations.
To achieve this, we optimize the interruption perturbation to maximize the distance between extracted features from the original and perturbed images:
@@
This loss ensures that the perturbation creates a consistent and significant distortion across the feature extractors of all targeted models. To achieve a more potent disruption than a simple MSE loss can provide, we introduce a feature enhancement loss $\mathcal{L}_{\mathrm{enh}}$ that amplifies discrepancies across three key levels: local, global and structural. This amplification is driven by a non-linear exponential function, which is critical for magnifying even small feature differences into a significant and destructive effect.
@@
Here, $\mathcal{F}_{\mathrm{enh}}$ denotes the enhanced feature. The term $w_{i}$ is a predefined weight vector that balances these three types of discrepancy. This amplification is regulated by a temperature scaling parameter $\sigma$. The term $\Delta$ represents a vector of local, global, and structural differences $(\mathrm{L_{D}},\mathrm{G_{D}},\mathrm{S_{D}})$, while $\hat{W}$ denotes the adaptive weights that control the enhancement strength. This component ensures that small feature differences are amplified into significant disruptions, resulting in a more comprehensive and destructive feature attack.
$\mathrm{L_{D}}=Norm({E_{i}(\hat{X})})-Norm({E_{i}(x)}),$ (4)
$G_{D}=\frac{U(E_{i}(\hat{X}))-U(E_{i}(x))}{S(E_{i}(x))+z},$ (5)
$\mathrm{S_{D}}=Att(Norm({E_{i}(\hat{X})}))-Att(Norm({E_{i}(x)})),$ (6)
@@
where $Norm(\cdot)$ denotes normalization function to standardize the feature map and $U(\cdot)$ calculates the mean of a feature map across its spatial dimensions. $S(\cdot)$ calculates the standard deviation of a feature map. $Att(\cdot)$ is an attention mechanism applied to normalized feature maps to compute and weight their structural feature. $z$ is a numerical stability constant.
@@
where $\gamma$ is the step size and $\epsilon$ is the perturbation budget. The projection operator ensures that the interruption perturbation remains within a bounded perturbation space.
@@ III-B Poisoning the Retraining Pipeline
While interruption is effective for real-time image distortion, long-term defense must prevent an attacker from using perturbed examples to fine-tune their model. Our approach creates an effective perturbation designed to disrupt face detection, which is the crucial first step in an attacker’s data processing pipeline.
@@
$L_{\text{feat}}=\sum_{i=1}^{N}w_{i}\cdot\|F_{\text{fused}}^{i}-(-F_{\text{fused}}^{i,\text{clean}}+\nu\cdot M_{d})\|^{2}.$ (12)
Here, $D_{i}(\text{x})$ is the feature extracted from the face detector $i_{th}$. $\alpha_{i}$ is the fusion weights that balance the contribution of each of the detectors and $w_{i}$ is the weight of the layer $i_{th}$. $N$ is the number of feature layers to attack. $\nu$ is a hyperparameter that controls the magnitude of the targeted feature-space margin. $F_{\text{fused}}^{i,\text{clean}}$ denotes the fused feature representation derived from the original image. The optimization of this poisoning loss is limited to specific regions $M_{d}$ defined by our intensity separation mechanism, and its technical details are given in the following section (III-C).
@@
$L_{total}=\mathcal{L}_{feat}-\mathcal{L}_{output}.$ (14)
@@ III-C Perturbation Fusion Optimization
The interruption perturbation and the subsequent poisoning perturbation are then synergistically combined into a unified perturbation. Directly combining the two defensive goals into a single weighted-sum loss function is often suboptimal, as competing gradients can lead to functional conflicts. To avoid this, our intensity separation mechanism decouples the two tasks spatially by utilizing the low-intensity regions of the primary interruption perturbation. This strategy allows both the interruption and poisoning functions to operate effectively without mutual interference, ensuring a more effective and efficient unified perturbation. The process of this mechanism is illustrated in Fig. 4.
@@
$\delta_{\text{final}}=\text{Clip}(W_{0}+\delta_{\text{t+1}},-\epsilon,\epsilon),$ (17)
@@
TABLE I: Comparison of Interruption Performance among CMUA, FOUND and TSDF. The avg is the average result of each indicator under this dataset. The best result is marked in bold.
@@
TABLE II: Comparison of Face Detector Performance Under Various Poisoning Attacks.
@@ IV-A Experimental Setup
Datasets: We conduct the experiments on the CelebA dataset [31] for training. The dataset comprises over 200,000 face images, from which we use the aligned and cropped version. To ensure a fair comparison with the SOTA baseline FOUND [20] while efficiently learning to disrupt the common and redundant features in facial images, we use a training set of 128 images. For evaluation, we then apply this generated perturbation to all images in the official test split. To further evaluate the generalization capability of TSDF, additional datasets are introduced, including the Labeled Faces in the Wild (LFW) dataset [32] and facial images extracted from the FaceForensics++ Original video dataset (FF++O) [33].
Deepfake Model: We evaluate the defense against four deepfake models, HiSD [34], AttentionGAN (AGGAN) [35], StarGAN [36], and AttGAN [37]. For StarGAN and AGGAN, we select 5 attributes: black hair, blonde hair, brown hair, gender, and age. In AttGAN’s training, the editing tasks are expanded to 14 attributes. Given HiSD’s advanced facial editing capabilities, the experiments incorporated modifications of the glasses attribute. To ensure experimental consistency and fairness, all deepfake models are trained on the CelebA dataset.
@@
Baseline: To rigorously evaluate our framework, our baselines were selected to represent the most foundational, representative and directly competing state-of-the-art (SOTA) methods. For interruption, we compare with the pioneering universal defense method CMUA-Watermark [15] and select FOUND [20] as a critical baseline. The latter’s principle of attacking an ensemble of feature extractors is similar to our own interruption module, making it the most relevant SOTA benchmark to validate our feature enhancement part. For the poisoning task, we benchmark against Face Poison [22], the most direct contemporary competitor that also targets face detection, and the foundational SSOD [23] to validate our effectiveness against detector-level attack strategies. Given the limited availability of public code, these poisoning baselines were rigorously re-implemented from their original papers to ensure a fair comparison.
Implementation Details: The hyperparameter configuration is aligned with existing baseline methods, while jointly optimizing the interruption and poisoning effects of the perturbations. The perturbation range for both attacks is set to (0.05, 0.05) to balance visual fidelity with attack efficacy. In the training stage, 50 and 20 iterations are performed for interruption and poisoning, as we observe in preliminary experiments that performance is saturated beyond these points. For the optimizer parameters in Equations 9 and 10, $\lambda$ and $\gamma$ are set to 0.1 and 0.001, respectively. The intensity separation threshold $\tau$ was set to 0.3, a value determined to provide the optimal balance between defensive functions, as validated in the ablation study.
@@ IV-C Compare the SOTA Methods
This section validates the effectiveness of the proposed TSDF framework through a series of direct comparisons with state-of-the-art (SOTA) methods.
@@
We evaluate the interruption capability of TSDF against SOTA methods, CMUA-Watermark [15] and FOUND [20], with detailed results presented in Table I. The results show that TSDF consistently delivers a more powerful and destructive effect on the final forged images across multiple datasets and models. A key advantage of the method is its superior ability to degrade the structural integrity of the forged output. This is most obvious in SSIM scores. On the CelebA dataset, TSDF achieves an average SSIM of 0.57, which is significantly lower than the scores of CMUA and FOUND, indicating superior disruption. This trend is particularly strong against models like StarGAN, where TSDF achieves a remarkable SSIM of just 0.24, far exceeding the disruptive effect of baseline. Furthermore, this disruptive force is reflected in other key metrics. On the LFW dataset, TSDF achieves the highest average L2mask of 0.28 and the highest average FID score of 282.32, indicating the most powerful disruption among the three methods. The minor difference in some of TSDF’s initial interruption metrics is an expected outcome of our dual-function design. Unlike single-function baselines that dedicate their entire budget to interruption, our framework must balance both interruption and poisoning capabilities within the same perturbation. Our ablation study quantitatively demonstrates that this balance is a deliberate design choice, made to achieve long-term persistence.
TABLE III: Degradation of Baseline Interruption Performance After Retraining.
@@
We then assess the effectiveness of TSDF’s poisoning component against face detectors, comparing it to random noise and reproduced versions of SSOD and FacePoison. The results in Table II are critical as they validate the core defense strategy. Instead of merely trying to survive an attacker’s adaptation, our framework fundamentally disrupts the retraining process. By significantly affecting the automated face detectors that an attacker relies on, our method makes large-scale and automated collection of training data impractical. This represents a fundamental shift in defensive strategy. The objective is no longer to create a perturbation that passively withstands retraining, but to design one that actively destroys the retraining process. This capability is crucial for achieving more persistent defense. For example, on the CelebA dataset, TSDF drastically reduces the F1-score for the RetinaFace detector to 0.78. This represents a far more successful attack compared to FacePoison and SSOD, which only achieve scores of 0.96 and 0.90 respectively. The advantage is even more pronounced against the S3FD detector, where our method achieves an F1-score of 0.72, while the other methods barely impact the detector’s performance. This superior performance is consistent across the LFW and FF++O datasets, where TSDF consistently achieves the lowest F1-score. These results validate that our framework’s poisoning component poses a significant threat to the face detection process. This capability is crucial for our two-stage defense, as it is the mechanism designed to prevent attackers from bypassing the interruption defense.
TABLE IV: Comparative Analysis of Defense Persistence After Retraining. The table compares the interruption effectiveness of CMUA, FOUND, and the TSDF before and after a simulated adaptive attack. A lower SSIM or a higher FID indicates a more effective and persistent interruption defense.
@@
TABLE V: Model Performance discrepancy After Retraining on TSDF Data. This table compares the performance of the original baseline model (Healthy) against the same model after it has been retrained on TSDF data (Poisoned).
@@
Figure 5: Visual comparison of the interruption effects. This figure compares the outputs of deepfake models when protected by CMUA, FOUND, and TSDF.
@@
TABLE VI: Comparison of SSIM and PSNR indicators on different datasets
@@
IV-D Analysis of Defense Persistence
This section is dedicated to the core contribution of this work, which is to validate the persistence of our TSDF framework. We conduct a direct comparative analysis that first establishes the failure of conventional SOTA defenses against an adaptive attack. We then demonstrate that our framework not only withstands the same attack but also enhances its defensive capabilities, thereby proving its superior persistence.
Baseline Performance under Adaptive Attack
To simulate a worst-case attack scenario, we create training datasets composed entirely of protected images. Specifically, we apply the SOTA interruption perturbations from CMUA and FOUND to the CelebA training set, creating two separate protected datasets. The deepfake models are then retrained on these respective datasets.
The results presented in Table III reveal a critical vulnerability in conventional interruption-only methods. For instance, with the CMUA method, the average SSIM score on the CelebA dataset increases from 0.64 to 0.79 after retraining, while the average FID score falls from 209.46 to 127.42. The rising SSIM indicates that the output of perturbed examples have become more similar to the original images, while the falling FID confirms they have become more realistic and less distorted. Both metrics prove that the interruption effect has been largely nullified. This degradation provides clear quantitative evidence that adaptive attacks severely compromise conventional interruption-only defenses, proving their lack of persistence. This highlights the key vulnerability our TSDF framework is designed to solve.
Comparative Analysis of TSDF’s Persistence
To provide a direct comparison against this established baseline of failure, we subjected our TSDF framework to the identical adaptive attack.
To validate the persistence of our framework, we conduct a direct comparative experiment against the SOTA interruption methods. To ensure a fair comparison, we evaluate all methods on the same two representative models, StarGAN and AttGAN. We simulated an adaptive attack by retraining these models on datasets protected by each respective defense. For TSDF, the retraining dataset was first processed by the poisoning mechanism. We then measured the interruption effectiveness of each defense before and after this adaptive attack to quantify their persistence.